20 x _______ = 5
How do you determine the unknown
fraction to solve the multiplication problem?

Found 2 solutions by MathLover1, math_tutor2020:
Answer by MathLover1(20850)
Answer by math_tutor2020(3821)


20 times something = 5
To isolate the "something", we need to divide both sides by 20. This undoes the multiplication.
Dividing by 20 is the same as multiplying by 1/20.

We end up with 5/20 = 1/4 that will go in the blank.

20 times 1/4 = 5